# FLAMINGO: combining kinetic SZ effect and galaxy-galaxy lensing measurements to gauge the impact of feedback on large-scale structure

### v1

@@ 4.2 The role of satellite galaxies
Here we use the FLAMINGO simulations and associated subhalo catalogs to explore the impact of the inclusion of satellite galaxies on the derived stacked lensing and kSZ effect profiles. In Fig. 4 we compare the ratios of the stacked lensing and kSZ signals for a central-only sample with that for the fiducial sample for different choices of the minimum stellar mass cut. For the lensing comparison (top panels) we show only the L1 and C2 bins for brevity, noting that the L2 and C1 bins give similar results. The solid curves show the ratios for the fiducial FLAMINGO feedback model and correspond to different minimum stellar masses: $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]$ values ranging between 10.9 and 11.6 in steps of 0.1.
@@
From our previous analysis, we found that minimum stellar masses in the range of $\approx 11.15$-$11.25$ best describe the selection for the BOSS LOWZ and CMASS samples. For these selections, Fig. 4 implies that the inclusion of satellites can boost the lensing and kSZ effect signals by up to $\approx 30\%$ and $\approx 20\%$, respectively. While these effects, which tend to be stronger for lower mass galaxies, are not large enough to alter the conclusion that stronger feedback is required to match the kSZ effect measurements relative to an X-ray-based calibration strategy, they nevertheless should be factored in for quantitative analyses. Furthermore, according to Fig. 4, as observations push to lower stellar masses, the role of satellites will become more significant and will need a careful accounting.
@@ 4.3 Implied halo masses
We note that if we apply the same minimum stellar mass cuts as in the fiducial selection but limit our analysis to central galaxies only, we find lower mean halo masses [^4] We find consistent constraints with the central galaxy only mean halo masses if, instead of computing the mean of the selected central galaxies, we simply fit a composite NFW profile + 2-halo term (using the halo bias model of Tinker et al. 2010) to the lensing data, using the Colossus package (Diemer, 2018). of $\log_{10}[M_{\textrm{500c}}/\textrm{M}_{\odot}]=13.32\pm 0.03$ and $13.13\pm 0.06$ for the LOWZ and CMASS samples, respectively. This confirms that the inclusion of satellite galaxies in the selection boosts the mean halo mass of the sample and by consequence also the stacked lensing and kSZ effect signals.
@@ 5.1.3 Unaccounted for systematics in the measurements
It is possible that, instead of there being a fundamental issue with the feedback in the simulations, there could be unaccounted for (or mischaracterised) systematic errors in the gas fraction, kSZ effect, or galaxy-galaxy lensing measurements. For example, there are considerable uncertainties in the X-ray selection function of galaxy groups (e.g., Pearson et al. 2017; Giles et al. 2022; Marini et al. 2024) and some studies make no attempt to account for selection effects. For the kSZ effect measurements, a significant underestimate of the bias in the velocity reconstruction, or from tSZ effect leakage in the kSZ effect stacking analysis, could potentially reconcile the measurements with the simulation predictions using X-ray-calibrated feedback. Ideally, we would use the same theory (hydrodynamical simulations) to select our systems, analyse them in precisely the same way as done for the real systems, and compare the processed observables in a like-with-like fashion. While we have taken important steps in this direction in the present study, the issues of the X-ray selection function and kSZ effect velocity reconstruction (and/or tSZ effect leakage) remain open questions. Finally, we have used galaxy-galaxy lensing measurements to constrain the selection of simulated galaxies and therefore any biases present in the lensing measurements will impact our feedback conclusions. However, as previously noted, we find excellent consistency in the derived mean halo masses from the lensing with that inferred from fitting to the large-scale projected clustering, suggesting any biases in the lensing masses are likely to be small.
@@ 5.2.1 Cosmic shear / 3D matter power spectrum
In Fig. 5 (left panel), we compare the predictions of the suppression of the matter power spectrum from fiducial FLAMINGO model and the strong fgas-$8\sigma$ variant with the recent constraints from Bigwood et al. (2024), Aricò et al. (2023), and Preston et al. (2023). The Preston et al. (2023) constraint comes from a joint analysis of the DES Y3 data and the Planck 2018 primary CMB. The constraints shown in Fig. 5 correspond to their best-fitting empirical ‘Amod’ model, which is intended to parameterise possible modifications of the non-linear part of the matter power spectrum due to baryonic physics and/or non-standard dark matter. The two constraints from Bigwood et al. (2024) come from an analysis of DES Y3 cosmic shear data (labelled WL) or from a joint analysis of the DES Y3 cosmic shear and the kSZ effect CMASS profile of Schaan et al. (2021) (labelled WL+kSZ), modelled using the BCEmu baryonification formalism (Schneider et al., 2019; Giri & Schneider, 2021). Note that Bigwood et al. (2024) do not adopt the small scale cuts employed in the fiducial DES Y3 cosmic shear analysis (e.g., Krause et al. 2021; Amon et al. 2022; Secco et al. 2022) that were designed to mitigate the impact of baryonic physics, but instead include the small scale measurements and use the baryonification formalism to model the baryonic effects, aided by the inclusion of kSZ effect data. Aricò et al. (2023) use the BACCO baryonification emulator [^5] We note that the BCEmu and BACCO model differences have been studied and shown to impact the shape of the predicted power spectrum (Grandis et al., 2024). A primary difference is that the former displaces formally all the particles in the simulation, whereas the latter displaces only particles in haloes. (Aricò et al., 2021) to model the DES Y3 cosmic shear data (including small scale measurements).
@@ 5.2.4 Summary
As the tSZ effect measurements appear strongly in tension with the standard model fit to the primary CMB and on the largest scales are generally insensitive to baryonic effects, it will be important to further scrutinise the robustness of these measurements, including the robustness of the component separation techniques employed to separate the tSZ effect from the other signals present in the data such as the CIB and radio sources. In this regard, cross correlations with other signals, such as cosmic shear, are invaluable since they will be differently affected by any residual contamination than is the case for the tSZ effect power spectrum. Relativistic effects should also be included in future analyses (e.g., Remazeilles et al. 2019; Remazeilles & Chluba 2024). Furthermore, we note that the baryonification formalism has recently been extended to tSZ effect modelling by Aricò & Angulo (2024) and it would be interesting to compare constraints from this methodology with those derived from the hydrodynamical simulations.
@@ 6 Conclusions
  • A simple minimum stellar mass-based selection employed on the simulations yields galaxy-galaxy lensing profiles which match the measurements of Amon et al. (2023) over the wide range of observed radii, with the best-fitting minimum stellar masses, but not the best-fitting haloes masses, depending slightly on the feedback model (Figs. 1 and 10) and the adopted cosmology. Higher minimum stellar masses (and halo masses) are required to match the LOWZ sample compared to the CMASS sample. Invoking more complicated selection functions that depend on both stellar mass and specific star formation rate yields results consistent with our fiducial analysis (see Appendix A). The derived mean halo masses for the CMASS and LOWZ samples are $\log_{10}[M_{\textrm{500c}}/\textrm{M}_{\odot}]=13.53\pm 0.02$ and $13.34\pm 0.04$ (at $2\sigma$ uncertainty), respectively.
  • Using the lensing measurements as a strong constraint on the selection for the kSZ effect predictions, we find that the fiducial FLAMINGO feedback model, which was calibrated on the X-ray gas fractions of low-redshift galaxy groups and clusters, is ruled out at the $7\sigma$ level by both the BOSS LOWZ and CMASS samples (Fig. 2). The strongest feedback variant considered in the FLAMINGO suite, the so-called fgas-$8\sigma$ model, provides a good match to the measurements, though is inconsistent with the local X-ray measurements.
@@ Acknowledgements
This work was supported by the Science and Technology Facilities Council (grant number ST/Y002733/1). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 769130). REA acknowledges support from the Spanish Ministry of Science under grant number PID2021-128338NB-I00. CSF acknowledges support from the European Research Council (ERC) through Advanced Investigator grant DMIDAS (GA 786910). This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
@@ Appendix A Galaxy selection including sSFR cuts
Appendix B Galaxy-galaxy lensing of additional FLAMINGO variations

### v2

@@ 4.2 The role of satellite galaxies
Here we use the FLAMINGO simulations and associated subhalo catalogs to explore the impact of the inclusion of satellite galaxies on the derived stacked lensing and kSZ effect profiles. In Fig. 4 we compare the ratios of the stacked lensing and kSZ signals for a central-only sample with that for the fiducial sample for different choices of the minimum stellar mass cut. For the lensing comparison (top panels) we show only the L1 and C2 bins for brevity, noting that the L2 and C1 bins give similar results. The solid curves show the ratios for the fiducial FLAMINGO feedback model and correspond to different minimum stellar masses: $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]$ values ranging between 10.9 and 11.6 in steps of 0.1 and in the legend we quote the satellite fractions corresponding to these minimum stellar mass selections.
@@
From our previous analysis, we found that minimum stellar masses in the range of $\approx 11.15$-$11.25$ best describe the selection for the BOSS LOWZ and CMASS samples. For these selections, Fig. 4 implies that the inclusion of satellites can boost the lensing and kSZ effect signals by up to $\approx 30\%$ and $\approx 20\%$, respectively. The effect is scale dependent, which is likely due to the mis-centring of satellites with respect to their hosts. One can therefore potentially constrain the satellite fraction through detailed measurements of the shape of the lensing and kSZ effect profiles.
While the effects of satellites are not large enough to alter our conclusion that stronger feedback is required to match the kSZ effect measurements relative to an X-ray-based calibration strategy, they nevertheless should be factored in for quantitative analyses. Furthermore, according to Fig. 4, as observations push to lower stellar masses, the role of satellites will become more significant (given the large satellite fractions) and will need a careful accounting.
@@ 4.3 Implied halo masses
We note that if we apply the same minimum stellar mass cuts as in the fiducial selection but limit our analysis to central galaxies only, we find lower mean halo masses [^5] We find consistent constraints with the central galaxy only mean halo masses if, instead of computing the mean of the selected central galaxies, we simply fit a composite NFW profile + 2-halo term (using the halo bias model of Tinker et al. 2010) to the lensing data, using the Colossus package (Diemer, 2018). of $\log_{10}[M_{\textrm{500c}}/\textrm{M}_{\odot}]=13.32\pm 0.03$ and $13.13\pm 0.06$ for the LOWZ and CMASS samples, respectively. This confirms that the inclusion of satellite galaxies in the selection boosts the mean halo mass of the sample and by consequence also the stacked lensing and kSZ effect signals. If we were instead to re-fit the minimum stellar mass for the central-only selection, the derived mean halo mass would increase, as expected. But note that a central-only selection does not provide a statistically good fit to the lensing data, so the derived mean halo mass would have questionable value.
@@ 5.1.3 Unaccounted for systematics in the measurements
It is possible that, instead of there being a fundamental issue with the feedback in the simulations, there could be unaccounted for (or mischaracterised) systematic errors in the gas fraction, kSZ effect, or galaxy-galaxy lensing measurements. For example, there are considerable uncertainties in the X-ray selection function of galaxy groups (e.g., Pearson et al. 2017; Giles et al. 2022; Marini et al. 2025) and some studies make no attempt to account for selection effects. For the kSZ effect measurements, a significant underestimate of the bias in the velocity reconstruction, or from tSZ effect leakage in the kSZ effect stacking analysis, could potentially reconcile the measurements with the simulation predictions using X-ray-calibrated feedback. Ideally, we would use the same theory (hydrodynamical simulations) to select our systems, analyse them in precisely the same way as done for the real systems, and compare the processed observables in a like-with-like fashion. While we have taken important steps in this direction in the present study, the issues of the X-ray selection function and kSZ effect velocity reconstruction (and/or tSZ effect leakage) remain open questions. Finally, we have used galaxy-galaxy lensing measurements to constrain the selection of simulated galaxies and therefore any biases present in the lensing measurements will impact our feedback conclusions. However, as previously noted, we find excellent consistency in the derived mean halo masses from the lensing with that inferred from fitting to the large-scale projected clustering, suggesting any biases in the lensing masses are likely to be small.
@@ 5.2.1 Cosmic shear / 3D matter power spectrum
In Fig. 5 (left panel), we compare the predictions of the suppression of the matter power spectrum from fiducial FLAMINGO model and the strong fgas-$8\sigma$ variant with the recent constraints from Bigwood et al. (2024), Aricò et al. (2023), and Preston et al. (2023) (see also Schaller et al. 2024a). The Preston et al. (2023) constraint comes from a joint analysis of the DES Y3 data and the Planck 2018 primary CMB. The constraints shown in Fig. 5 correspond to their best-fitting empirical ‘Amod’ model, which is intended to parameterise possible modifications of the non-linear part of the matter power spectrum due to baryonic physics and/or non-standard dark matter. The two constraints from Bigwood et al. (2024) come from an analysis of DES Y3 cosmic shear data (labelled WL) or from a joint analysis of the DES Y3 cosmic shear and the kSZ effect CMASS profile of Schaan et al. (2021) (labelled WL+kSZ), modelled using the BCEmu baryonification formalism (Schneider et al., 2019; Giri & Schneider, 2021). Note that Bigwood et al. (2024) do not adopt the small scale cuts employed in the fiducial DES Y3 cosmic shear analysis (e.g., Krause et al. 2021; Amon et al. 2022; Secco et al. 2022) that were designed to mitigate the impact of baryonic physics, but instead include the small scale measurements and use the baryonification formalism to model the baryonic effects, aided by the inclusion of kSZ effect data. Aricò et al. (2023) use the BACCO baryonification emulator [^6] We note that the BCEmu and BACCO model differences have been studied and shown to impact the shape of the predicted power spectrum (Grandis et al., 2024). A primary difference is that the former displaces formally all the particles in the simulation, whereas the latter displaces only particles in haloes. (Aricò et al., 2021) to model the DES Y3 cosmic shear data (including small scale measurements).
@@ 5.2.4 Summary
As the tSZ effect measurements appear strongly in tension with the standard model fit to the primary CMB and on the largest scales are generally insensitive to baryonic effects, it will be important to further scrutinise the robustness of these measurements, including the robustness of the component separation techniques employed to separate the tSZ effect from the other signals present in the data such as the CIB and radio sources. In this regard, cross correlations with other signals, such as cosmic shear, are invaluable since they will be differently affected by any residual contamination than is the case for the tSZ effect power spectrum. Relativistic effects should also be included in future analyses (e.g., Remazeilles et al. 2019; Remazeilles & Chluba 2025). Furthermore, we note that the baryonification formalism has recently been extended to tSZ effect modelling by Aricò & Angulo (2024) and it would be interesting to compare constraints from this methodology with those derived from the hydrodynamical simulations.
@@ 6 Conclusions
  • A simple minimum stellar mass-based selection employed on the simulations yields galaxy-galaxy lensing profiles which match the measurements of Amon et al. (2023) over the wide range of observed radii, with the best-fitting minimum stellar masses, but not the best-fitting haloes masses, depending slightly on the feedback model (Figs. 1 and 11) and the adopted cosmology. Higher minimum stellar masses (and halo masses) are required to match the LOWZ sample compared to the CMASS sample. Invoking more complicated selection functions that depend on both stellar mass and specific star formation rate yields results consistent with our fiducial analysis (see Appendix A). The derived mean halo masses for the CMASS and LOWZ samples are $\log_{10}[M_{\textrm{500c}}/\textrm{M}_{\odot}]=13.53\pm 0.02$ and $13.34\pm 0.04$ (at $2\sigma$ uncertainty), respectively.
  • Using the lensing measurements as a strong constraint on the selection for the kSZ effect predictions, we find that the fiducial FLAMINGO feedback model, which was calibrated on the X-ray gas fractions of low-redshift galaxy groups and clusters, is ruled out at the $5\sigma$ level by both the BOSS LOWZ and CMASS samples (Fig. 2). The strongest feedback variant considered in the FLAMINGO suite, the so-called fgas-$8\sigma$ model, provides a good match to the measurements, though is inconsistent with the local X-ray measurements.
@@ Acknowledgements
We thank the referee for a helpful report that improved the paper. This work was supported by the Science and Technology Facilities Council (grant number ST/Y002733/1). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 769130). REA acknowledges support from the Spanish Ministry of Science under grant number PID2021-128338NB-I00. CSF acknowledges support from the European Research Council (ERC) through Advanced Investigator grant DMIDAS (GA 786910). This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
@@ Appendix A Galaxy selection including sSFR cuts
Appendix B Resolution dependence
In Fig. 10 we compare the predictions of the high-resolution calibrated FLAMINGO model (L1_m8) with the fidicual resolution calibrated model (L1_m9). As per our default analysis, to predict the kSZ effect we first determine the minimum galaxy stellar mass limit that best matches the galaxy-galaxy lensing measurements of Amon et al. (2023) independently for the two models. For the high-resolution model, the best-fit masses are $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.35\pm 0.04$ and $11.26\pm 0.06$ for the LOWZ and CMASS samples, respectively. These are the best-fit masses from fitting the two redshift bins in each sample (e.g., LOWZ-L1 and LOWZ-L2) jointly. These values are somewhat higher than the best-fit joint masses for the fiducial resolution run, which are $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.26\pm 0.02$ and $11.18\pm 0.04$ for LOWZ and CMASS, respectively, (though only the LOWZ difference is statistically significant). This difference is expected since, although both models were independently calibrated to the observed galaxy stellar mass function, the calibration to the data is not perfect. Furthermore, the simulations were only calibrated up to $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.5$ and higher mass do affect the sample. We note that in the galaxy stellar mass function the high-resolution run is offset from the fiducial resolution run (and the observational measurements) by $\approx 0.1-0.2$ dex at a stellar mass scale of $\sim 10^{11}\textrm{M}_{\odot}$, resulting in larger stellar masses compared to the fiducial resolution run at a halo mass scale of $\sim 10^{13}\textrm{M}_{\odot}$ (see fig. 9 of Schaye et al. 2023). Thus, to match the galaxy-galaxy lensing signal, a higher stellar mass cut is required for the high-resolution run.
When the galaxy selection is fixed by the lensing, the predicted kSZ effect profiles are similar (see Fig. 10), typically deviating from each by less than the observational measurement errors, and both are in strong tension with the measurements of Schaan et al. (2021).
Figure 10: As Fig. 2 but comparing the fiducial resolution calibrated model (L1_m9) with the high-resolution calibrated model (L1_m8). Overall the profiles are similar (typically deviating from each by less than the observational measurement errors) and both are in strong tension with the measurements of Schaan et al. (2021).
Appendix C Galaxy-galaxy lensing of additional FLAMINGO variations